# Inferring Outcome Means of Exponential Family Distributions Estimated by Deep Neural Networks

### v2

@@ Abstract
Despite the widespread use of deep neural networks (DNNs) for prediction, inference on estimated means for categorical or exponential family outcomes remains underexplored. We address this gap by framing the problem within the generalized linear models (GLMs) framework and developing a rigorous statistical approach for inference on DNN-estimated means. To address a key limitation of assuming independence between prediction errors and input variables in the literature, which often fails in GLMs, we introduce a truncation technique that partitions the problem into regimes with distinct noise behaviors, enabling refined analysis and robust theoretical guarantees under general GLM frameworks. To implement inference, we consider an Ensemble Subsampling Method (ESM) that leverages U-statistics and the Hoeffding decomposition to construct reliable confidence intervals. This method enables model-free variance estimation and accounts for heterogeneity among individuals in the population. Through extensive simulations across Binary, Poisson and Binomial models, we demonstrate the effectiveness and efficiency of our method. We further apply the method to real-world data from the eICU dataset to predict patient readmission risks, providing actionable insights for clinical decision-making.
@@ 1 Introduction
In supervised learning (Hastie et al., 2009), the focus is often on the estimation of the conditional expectation $\mathbb{E}[y|\mathbf{x}]$, which represents the optimal prediction of the target variable $y$ given covariates $\mathbf{x}$. The notion is fundamental because for a pair $(\mathbf{x},y)$, the conditional expectation satisfies:
$\mathbb{E}[\ell(y,y(\mathbf{x}))]\geq\mathbb{E}[\ell(y,\mathbb{E}[y|\mathbf{x}])]$
for any predictor $y(\mathbf{x})$, where the loss function $\ell(y,y(\mathbf{x}))$ measures the discrepancy between the true label $y$ and the prediction $y(\mathbf{x})$. Numerous methods have been developed to estimate $\mathbb{E}[y|\mathbf{x}]$, including linear, parametric and nonparametric regression, kernel regression (Nadaraya, 1964), support vector regression (Drucker et al., 1996), and deep neural networks (LeCun et al., 2015).
@@
Deep neural networks (DNNs) present remarkable capabilities in modeling $\mathbb{E}[y|\mathbf{x}]$, and often outperform traditional techniques in accuracy and scalability. However, the statistical inference surrounding the estimation of $\mathbb{E}[y|\mathbf{x}]$, especially when $y$ is categorical or belongs to the exponential family, remains underdeveloped. We aim to address this gap by establishing a rigorous framework for statistical inference on $\mathbb{E}[y|\mathbf{x}]$ estimated by DNNs under a generalized nonparametric regression model (GNRM) setting [see Eq. (2.3)].
Modern inference approaches include distribution-free conformal prediction inference (Lei et al., 2018; Angelopoulos and Bates, 2021; Huang et al., 2024), extensions of jackknife inference within conformal prediction (Alaa and Van Der Schaar, 2020; Kim et al., 2020; Barber et al., 2021), debiasing methods (Athey et al., 2018; Guo et al., 2021), residual-based bootstrap techniques (Zhang and Politis, 2023), and the use of $U$ statistics in regressions, random forests and neural networks (Mentch and Hooker, 2016; Wager and Athey, 2018; Schupbach et al., 2020; Wang and Wei, 2022; Fei and Li, 2024). Several limitations remain. Conformal prediction, for example, provides prediction intervals under continuous regression settings by accounting for the randomness of both $y$ and $\mathbb{E}[y|\mathbf{x}]$. However, the resulting intervals tend to be conservative, producing wide bounds and leading to reduced efficiency compared to methods that directly model the uncertainty of $\mathbb{E}[y|\mathbf{x}]$. Debiasing techniques such as those in (Athey et al., 2018; Guo et al., 2021) yield useful intervals, but are restricted to models with linear functions of $\mathbf{x}$, limiting their applicability to DNNs. Similarly, the residual bootstrap method proposed in (Zhang and Politis, 2023) is adapted to linear regression frameworks and does not readily extend to more complex prediction scenarios. Mentch and Hooker (2016) and Wager and Athey (2018) provided insights into quantifying prediction uncertainty for random forests, and Schupbach et al. (2020) explored U-statistics in neural networks, albeit with limited theoretical justification. Recently, Fei and Li (2024) examined U-statistics in neural networks rigorously, but their results remain tied to continuous outcomes, leaving open questions of whether these methods can be applied to GNRMs such as nonparametric logistic and Poisson regressions.
@@
Theoretically, for valid inference on $\mathbb{E}[y|\mathbf{x}]$, we need to establish the convergence rate of neural networks in a GNRM framework. Since the universal approximation theorem (Hornik et al., 1989), there has been substantial progress in understanding the convergence rates of neural networks; see McCaffrey and Gallant (1994); Kohler and Krzyżak (2005); Yarotsky (2017, 2018); Shen et al. (2019); Kohler and Langer (2021); Lu et al. (2021); Shen et al. (2021); Jiao et al. (2023); Yan and Yao (2023); Fan et al. (2024); Wang et al. (2024); Bhattacharya et al. (2024). For instance, Shen et al. (2021) established the convergence rate under nonparametric regression with the assumption $p$-th moment of response is bounded for some $p>1$; Fan et al. (2024) explored nonasymptotic error bounds of the estimators under several loss functions, such as the least-absolute deviation loss and Huber loss; Wang et al. (2024) proposed an efficient and robust nonparametric regression estimator based on an estimated maximum likelihood estimator; Fan and Gu (2024); Bhattacharya et al. (2024) investigated deep neural network estimators for nonparametric interaction models with diverging input dimensions, focusing on theoretical properties such as minimax optimality, approximation error control, and sparsity-driven regularization. However, quantifying uncertainty of DNN-estimated $\mathbb{E}[y|\mathbf{x}]$ within the framework of GNRM introduces additional challenges. The existing approaches typically assume a nonparametric regression framework, where the distribution of $y-\mathbb{E}[y|\mathbf{x}]$ is independent of the covariates $\mathbf{x}$. This assumption simplifies the theoretical analysis and the derivation of convergence results (Takezawa, 2005; Györfi et al., 2006; Fan, 2018). In generalized nonparametric regression settings, however, the assumption often fails as the distribution of $y-\mathbb{E}[y|\mathbf{x}]$ may still depend on $\mathbf{x}$. Consequently, the standard techniques for bounding error terms and controlling variance no longer apply, and the established results on uniform convergence or concentration inequalities (Bartlett et al., 2020) are invalid. As a result, it is unclear whether these methods are applicable for drawing inference on GNRMs estimated by DNNs.
Our contributions in addressing these challenges are as follows. First, we propose a DNN estimator for subject-specific means under GNRMs and develop a rigorous inference framework. To establish the convergence rate — crucial for demonstrating the feasibility of using the estimator for both estimating $f_{0}$ and conducting inference — we account for the dependence of the distribution of $y-\mathbb{E}[y|\mathbf{x}]$ on $\mathbf{x}$, which arises from heteroskedasticity in GNRMs. We apply truncations to partition the analysis into bounded and unbounded noise regimes. Standard convergence results hold in the bounded regime, while in the unbounded regime, integral bounds and concentration inequalities ensure that extreme noise contributions are asymptotically negligible. This yields rigorous error bounds under general GNRM conditions. As opposed to Schmidt-Hieber (2020) that focused on continuous outcomes, our work accommodates a broader class of outcomes, both continuous and discrete, encompassing the work of (Schmidt-Hieber, 2020) as a special case. This demonstrates that similar principles can be extended to the wider GNRM framework. Second, to quantify the uncertainty of the estimates, we propose an ensemble subsampling method (ESM) under GNRMs, developed from U-statistics (Frees, 1989; Hoeffding, 1992; Lee, 2019; Boroskikh, 2020) and Hoeffding decomposition (Hoeffding, 1992). Our model-free variance estimator leverages the intrinsic properties of Hoeffding decomposition and accommodates varying population levels as well as individual variance structures.
@@
In Section 2, we introduce deep neural networks (DNNs), generalized nonparametric regression models (GNRMs) and the ensemble subsampling method (ESM). Section 3 presents theoretical results for the convergence rate of DNNs under the GNRM framework, as well as results for statistical inference using ESM. In Section 4, we perform simulations to evaluate the finite sample performance of our method. Section 5 applies the methodology to an electronic Intensive Care Unit (eICU) dataset to predict the number of ICU readmissions. The eICU dataset is a large-scale, multi-center dataset containing anonymized health data from ICU patients, as part of a collaboration between the Massachusetts Institute of Technology and several healthcare organizations. The results may inform identification of high-risk individuals, enabling caregivers to offer personalized care. Section 6 concludes the paper with discussions. The proofs of the main theorems are given in Appendix A, while additional numerical experiments and technical proofs are included in the Supplementary materials.
@@ 2 The Preamble
We employ lower letters (e.g. $a$) for scalars and boldface letters for vectors and matrices (e.g. $\mathbf{a}$ and $\mathbf{A}$). For any matrix $\mathbf{A}$, we use $\|\mathbf{A}\|_{0}$ and $\|\mathbf{A}\|_{\infty}$ to denote its zero norm and infinity norm, respectively. The set of natural and real numbers are denoted by $\mathbb{N}$ and $\mathbb{R}$ respectively. For any function $f$ mapping a vector to $\mathbb{R}$, we use $\|f\|_{\infty}$ to denote its infinity norm. For an integer $n$, $[n]=\{1,\ldots,n\}$. We write $X_{1}(n)=O(X_{2}(n))$ or $X_{1}(n)\precsim X_{2}(n)$ if for any $\varepsilon>0$, there exists $C>0$ such that $\mathbb{P}(|X_{1}(n)/X_{2}(n)|>C)\leq\varepsilon$ for all $n$. We denote $X_{1}(n)=o(X_{2}(n))$ if $\{X_{1}(n)/X_{2}(n)\}$ converges to 00 in probability, and $X_{1}(n)\sim X_{2}(n)$ if $\{X_{1}(n)/X_{2}(n)\}$ converges to $1$ in probability. When clear from context, we use the same symbol for a random variable and its realized value to avoid redundancy.
@@ 2.1 Deep Neural Networks
Let $L$ be a positive integer and $\mathbf{p}=(p_{0},...,p_{L},p_{L+1})^{\top}$ be a vector of positive integers. An $(L+1)$-layer DNN with layer-width $\mathbf{p}$ is defined as
@@
with the recursive expression
$f_{l}(\mathbf{x})=\sigma(\mathbf{W}_{l-1}f_{l-1}(\mathbf{x})+\mathbf{v}_{l-1}),\quad f_{1}(\mathbf{x})=\sigma(\mathbf{W}_{0}\mathbf{x}+\mathbf{v}_{0}).$
Here, $\mathbf{W}_{l}\in\mathbb{R}^{p_{l+1}\times l}$ and vectors $\mathbf{v}_{l}\in\mathbb{R}^{p_{l+1}}$ ($l\in[L]$) are the parameters of this DNN, and $\sigma:\mathbb{R}\to\mathbb{R}$, an activation function, is applied element-wise to vectors. A commonly used activation function is the rectified linear unit (ReLU) function (Nair and Hinton, 2010), or $\sigma(x)=\max(x,0).$ We focus on it due to its piecewise linearity and projection properties, with potential extensions to other activation functions as discussed in Section 6. In (2.1), $L$ denotes the depth of the neural network, $\mathbf{p}$ specifies the width of each layer, with $p_{0}$ corresponding to the input dimension, $p_{1},\dots,p_{L}$ the dimensions of the $L$ hidden layers, and $p_{L+1}$ the dimension of the output layer.
Fully connected DNNs have a large number of parameters, which can lead to overfitting by making the network memorize training data rather than generalize. Introducing sparsity and boundedness reduces model complexity and improves generalization. We focus on the following class of $s$-sparse and $F$-bounded networks for our theoretical analysis:
@@ 2.2 Generalized Nonparametric Regression
We consider outcomes from the exponential family, encompassing widely used distributions such as the normal, Bernoulli, Poisson, and binomial. For a random variable $y$ belonging to the exponential family, its probability density (or mass) function is given by
$p(y|\theta)=h(y)\exp\left\{\theta y-\psi(\theta)\right\},$
where $\theta$ is the natural parameter, $h(y)$ is the base measure, and $\psi(\theta)$ is the log-partition function to ensure normalization. Following Fei and Li (2021), we have omitted the dispersion parameter for simplicity, as it does not impact the estimation of the means in our cases.
@@
where $f_{0}:\mathbb{R}^{d}\to\mathbb{R}$ is the nonparametric function linking $\mathbf{x}$ to $\theta$. Model (2.3) is referred to as a Generalized Nonparametric Regression Model (GNRM), where the outcomes follow a distribution from the exponential family. To predict $y_{{\mathop{\mathrm{new}}}}$ given a new test point $\mathbf{x}_{{\mathop{\mathrm{new}}}}$, a natural choice is $\mathbb{E}(y_{{\mathop{\mathrm{new}}}}|\mathbf{x}_{{\mathop{\mathrm{new}}}})=\psi^{\prime}(f_{0}(\mathbf{x}_{{\mathop{\mathrm{new}}}}))$, where the inverse of function $\psi^{\prime}(\cdot)$ is commonly known as the link function. Consequently, it is of interest to estimate $f_{0}$. By the universal approximation theorem (Hornik et al., 1989), deep neural networks are suitable candidates for estimating $f_{0}$. We propose a DNN estimator, $\widehat{f}^{*}_{n}$, which minimizes the negative log likelihood function (2.3) over the class of DNNs defined in (2.2), that is,
$\widehat{f}^{*}_{n}=\operatorname{argmin}_{f\in\mathcal{F}(L,\mathbf{p},s,F)}\frac{1}{n}\sum_{i=1}^{n}\{-y_{i}f(\mathbf{x}_{i})+\psi(f(\mathbf{x}_{i}))\}.$ (2.4)
@@
Nonparametric Gaussian Regression: Given $\mathbf{x}$, we assume that the response variable follows a Gaussian distribution with mean $f_{0}(\mathbf{x})$ and variance $\sigma^{2}$. In this setting, the variance parameter $\sigma^{2}$ does not affect the estimation of the mean function $f_{0}(\mathbf{x})$, which can be seen by examining the likelihood and optimization. The likelihood function for a nonparametric Gaussian regression model is given by $p(y|\mathbf{x})=\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(y-f_{0}(\mathbf{x}))^{2}}{2\sigma^{2}}}$. An application of (2.4) to this setting yields:
$\widehat{f}^{*}_{n}$ $=$ $\operatorname{argmin}_{f\in\mathcal{F}(L,\mathbf{p},s,F)}\frac{1}{n}\sum_{i=1}^{n}\{-y_{i}f(\mathbf{x}_{i})+\frac{1}{2}f^{2}(\mathbf{x}_{i})\}$
@@
which corresponds to minimizing the empirical mean squared error (MSE), as considered by Schmidt-Hieber (2020).
Nonparametric Logistic Regression: To model binary outcomes, we set $h(y)=1$, $\psi(\theta)=\log(1+e^{\theta})$ in (2.3), and let $y\in\{0,1\}$. The estimator (2.4) in this setting can be written as:
$\widehat{f}^{*}_{n}=\operatorname{argmin}_{f\in\mathcal{F}(L,\mathbf{p},s,F)}\frac{1}{n}\sum_{i=1}^{n}(-y_{i}f(\mathbf{x}_{i})+\log[1+\exp\{f(\mathbf{x}_{i})\}]).$
@@
$\widehat{f}^{*}_{n}=\operatorname{argmin}_{f\in\mathcal{F}(L,\mathbf{p},s,F)}\frac{1}{n}\sum_{i=1}^{n}[\exp\{f(\mathbf{x}_{i})\}-y_{i}f(\mathbf{x}_{i})].$
@@
Nonparametric Binomial Regression: Consider that, given covariates $\mathbf{x}$, the response variable $y$ follows a Binomial distribution with a fixed number of trials $n_{\text{trial}}$ and the success probability $p(\mathbf{x})=\frac{1}{1+e^{-f_{0}(\mathbf{x})}}$. This corresponds to $h(y)={n_{\text{trial}}\choose{y}}$ and $\psi(\theta)=n_{\text{trial}}\log(1+e^{\theta})$. In this setting, the estimator (2.4) is specified as

### v3

# Inference for Deep Neural Network Estimators in Generalized Nonparametric Models

@@ Abstract
While deep neural networks (DNNs) are used for prediction, inference on DNN-estimated subject-specific means for categorical or exponential family outcomes remains underexplored. We address this by proposing a DNN estimator under generalized nonparametric regression models (GNRMs) and developing a rigorous inference framework. Unlike existing approaches that assume independence between estimation errors and inputs to establish the error bound, a condition often violated in GNRMs, we allow for dependence and our theoretical analysis demonstrates the feasibility of drawing inference under GNRMs. To implement inference, we consider an Ensemble Subsampling Method (ESM) that leverages U-statistics and the Hoeffding decomposition to construct reliable confidence intervals for DNN estimates. We show that, under GNRM settings, ESM enables model-free variance estimation and accounts for heterogeneity among individuals in the population. Through simulations under nonparametric logistic, Poisson, and binomial regression models, we demonstrate the effectiveness and efficiency of our method. We further apply the method to the electronic Intensive Care Unit (eICU) dataset, a large scale collection of anonymized health records from ICU patients, to predict ICU readmission risk and offer patient-centric insights for clinical decision making.
Keywords: Deep Neural Networks, Inference, Generalized Nonparametric Regression Models, Ensemble Subsampling Method
@@ 1 Introduction
Given an outcome variable $y$ and covariates $\mathbf{x}$, estimating the conditional expectation $\mathbb{E}(y|\mathbf{x})$ is central to supervised learning. Under squared loss, $\mathbb{E}(y|\mathbf{x})$ is the optimal estimator that minimizes the expected estimation error among all measurable functions, making it a natural and widely adopted target in statistical modeling and machine learning. Even when optimality does not hold under more general losses, $\mathbb{E}(y|\mathbf{x})$ remains a valuable target, offering insight into the relationship between covariates and outcomes. It serves as a foundational component for constructing more complex quantities across domains such as conditional variances, counterfactual means, and risk-adjusted estimates (Hastie et al., 2009). Methods that estimate $\mathbb{E}(y|\mathbf{x})$ include linear, parametric and nonparametric regression, kernel regression (Nadaraya, 1964), support vector regression (Drucker et al., 1996), random forests (Breiman, 2001) and deep neural networks (DNNs) (LeCun et al., 2015).
@@
With the rapid advancement of artificial intelligence, DNN–based estimators have gained widespread popularity. However, the statistical inference surrounding the DNN estimator of $\mathbb{E}(y|\mathbf{x})$, especially when $y$ is categorical or belongs to the exponential family, remains underdeveloped. We address this gap by establishing a rigorous framework for inference on $\mathbb{E}(y|\mathbf{x})$ estimated by DNNs under a generalized nonparametric regression model (GNRM) setting [see Eq. (2.3)].
Modern inference approaches include distribution-free conformal estimation inference (Lei et al., 2018; Angelopoulos & Bates, 2021; Huang et al., 2024), extensions of jackknife inference within conformal estimation (Alaa & Van Der Schaar, 2020; Kim et al., 2020; Barber et al., 2021), debiasing methods (Athey et al., 2018; Guo et al., 2021), residual-based bootstrap techniques (Zhang & Politis, 2023), and the use of $U$ statistics in regressions, random forests and neural networks (Mentch & Hooker, 2016; Wager & Athey, 2018; Schupbach et al., 2020; Wang & Wei, 2022; Fei & Li, 2024). Several distinctions and limitations remain. For example, conformal prediction constructs distribution-free prediction sets for $y$, targeting outcome coverage rather than inference on $\mathbb{E}(y|\mathbf{x})$. Its intervals thus absorb both outcome noise and model uncertainty, making them conservative and inefficient for mean inference. Clinically, inference on risk classifications derived from patients’ risk profiles, which are based on the estimated conditional mean (i.e., risk score), is more actionable, as medical decisions (e.g., triage or treatment intensity) rely on risk stratifications rather than prediction sets (Lundberg et al., 2018; Van Calster et al., 2019). Jackknife and related leave-one-out methods assume stable, nearly independent residuals, conditions violated by DNNs whose estimates depend heavily on individual samples. Debiasing approaches (Athey et al., 2018; Guo et al., 2021) and residual bootstraps (Zhang & Politis, 2023) remain limited to linear models, while resampling-based inference for high-dimensional GLMs (Fei & Li, 2021) is fully parametric. Recent work on uncertainty quantification for random forests (Mentch & Hooker, 2016; Wager & Athey, 2018) and neural networks (Schupbach et al., 2020; Fei & Li, 2024) offers valuable insights but remains largely confined to continuous outcomes, leaving inference for GNRMs such as logistic and Poisson regressions unresolved.
@@
Theoretically, for valid inference on $\mathbb{E}(y|\mathbf{x})$, we need to establish the convergence rate of neural networks in a GNRM framework. Since the universal approximation theorem (Hornik et al., 1989), there has been much progress in understanding the convergence rates of neural networks (McCaffrey & Gallant, 1994; Kohler & Krzyżak, 2005; Yarotsky, 2017; Shen et al., 2019; Kohler & Langer, 2021; Lu et al., 2021; Shen et al., 2021; Jiao et al., 2023; Fan et al., 2024; Wang et al., 2024; Bhattacharya et al., 2024; Yan et al., 2025). For instance, Shen et al. (2021) established the convergence rate under nonparametric regression assuming the $p$-th moment of response is bounded for some $p>1$; Fan et al. (2024) explored nonasymptotic error bounds of the estimators under several loss functions, such as the least-absolute deviation loss and Huber loss; Wang et al. (2024) proposed an efficient and robust nonparametric regression estimator based on an estimated maximum likelihood estimator; Fan & Gu (2024); Bhattacharya et al. (2024) investigated deep neural network estimators for nonparametric interaction models with diverging input dimensions, focusing on theoretical properties such as minimax optimality, approximation error control, and sparsity-driven regularization. However, quantifying uncertainty of DNN-estimated $\mathbb{E}(y|\mathbf{x})$ within the framework of GNRM introduces additional challenges. The existing approaches typically assume a nonparametric regression framework, where the distribution of $y-\mathbb{E}(y|\mathbf{x})$ does not depend on $\mathbf{x}$. This assumption simplifies the theoretical analysis and the derivation of convergence results (Takezawa, 2005; Györfi et al., 2006; Fan, 2018). In generalized nonparametric regression settings, however, the assumption often fails due to heteroskedasticity, as the distribution of $y-\mathbb{E}(y|\mathbf{x})$ may still depend on $\mathbf{x}$. Consequently, the standard techniques for bounding error terms and controlling variance no longer apply, and the established results on uniform convergence or concentration inequalities (Bartlett et al., 2020) are invalid. As a result, it is unclear whether these methods are applicable for drawing inference on GNRMs estimated by DNNs.
Our methodological and theoretical contributions are as follows. First, we propose a DNN estimator for subject specific means under GNRMs and develop an inference framework that explicitly accounts for covariate-dependent error distributions induced by heteroskedasticity. We provide new proof techniques, including new criterion loss function, covering-number based concentration and a truncation strategy that handles exponential-tailed, heteroskedastic noise, improving the results of Schmidt-Hieber (2020). For example, with the general loss induced by GNRMs, the quadratic-based arguments of Schmidt-Hieber (2020) no longer apply. This necessitates a complete redevelopment of the error bounds, derived instead from the intrinsic curvature and Lipschitz properties in the new GNRM setting. As a result, all estimation error analyses are novel and move beyond Gaussian approximations. Additionally, the presence of exponential tailed, heteroskedastic noise complicates uniform convergence, which we address through a truncation strategy. The truncation scheme partitions the analysis into bounded and unbounded noise regimes, where standard convergence arguments apply in the former and concentration inequalities control the latter, yielding sharper error bounds. These theoretical tools enable us to extend the continuous outcome setting of Schmidt-Hieber (2020), accommodating both continuous and discrete outcomes, and thereby generalizing their framework to the broader GNRM class. Second, to quantify the uncertainty of the estimates, we propose an ensemble subsampling method (ESM) under GNRMs, developed from U-statistics (Frees, 1989; Hoeffding, 1992; Lee, 2019; Boroskikh, 2020) and Hoeffding decomposition (Hoeffding, 1992). Our model-free variance estimator leverages the intrinsic properties of Hoeffding decomposition and accommodates varying population levels as well as individual variance structures. Moreover, we employ incomplete U-statistics to reduce computational cost, enabling the resample size to grow proportionally with the sample size while limiting the total number of resamples. To mitigate the resulting Monte Carlo bias, we refine the variance estimator of Wager & Athey (2018) for bias correction.
@@
In Section 2, we introduce deep neural networks (DNNs), generalized nonparametric regression models (GNRMs) and the ensemble subsampling method (ESM). Section 3 presents theoretical results for the convergence rate of DNNs under the GNRM framework, as well as results for statistical inference using ESM. In Section 4, we perform simulations to evaluate the finite sample performance of our method. Section 5 demonstrates the proposed methodology using the electronic Intensive Care Unit (eICU) dataset to estimate both the probability and expected number of ICU readmissions based on patient characteristics, with accompanying statistical inference. The eICU dataset is a multi-center resource comprising anonymized clinical data from ICU patients, developed through a collaboration between the Massachusetts Institute of Technology and multiple healthcare partners. The findings may support the identification of high-risk individuals and guide the delivery of personalized, risk-informed care. Section 6 concludes the paper with discussions. Additional experiments and proofs are provided in the Supplementary Material.
Notation. We employ lower letters (e.g. $a$) for scalars and boldface letters for vectors and matrices (e.g. $\mathbf{a}$ and $\mathbf{A}$). For any matrix $\mathbf{A}$, we use $\|\mathbf{A}\|_{0}$ and $\|\mathbf{A}\|_{\infty}$ to denote its zero norm and infinity norm, respectively. The set of natural and real numbers are denoted by $\mathbb{N}$ and $\mathbb{R}$ respectively. For any function $f$ mapping a vector to $\mathbb{R}$, we use $\|f\|_{\infty}$ to denote its infinity norm. For an integer $n$, $n]=\{1,\ldots,n\}$. We write $X_{1}(n)=O(X_{2}(n))$ or $X_{1}(n)\precsim X_{2}(n)$ or $X_{2}(n)\succsim X_{1}(n)$ if there exist $C>0$ and $N_{0}>0$ such that $|X_{1}(n)|\leq C|X_{2}(n)|$ when $n>N_{0}$. We denote $X_{2}(n)\asymp X_{1}(n)$ if $X_{1}(n)=O(X_{2}(n))$ and $X_{2}(n)=O(X_{1}(n))$. We denote $X_{1}(n)=o(X_{2}(n))$ if $X_{1}(n)/X_{2}(n)\to 0$, and $X_{1}(n)\sim X_{2}(n)$ if $\{X_{1}(n)/X_{2}(n)\}$ converges to $1$. When clear from context, we add the index $p$ (for probability) such as $O_{p}$ for a random variable and its realized value to avoid redundancy.
@@ 2.1 Deep Neural Networks
Let $L$ be a positive integer denoting the depth of a neural network, and $\mathbf{p}=(p_{0},...,p_{L},p_{L+1})^{\top}$ be a vector of positive integers specifying the width of each layer, with $p_{0}$ corresponding to the input dimension, $p_{1},\dots,p_{L}$ the dimensions of the $L$ hidden layers, and $p_{L+1}$ the dimension of the output layer. An $(L+1)$-layer DNN with layer-width $\mathbf{p}$ is defined as
@@
with the recursive expression $f_{l}(\mathbf{x})=\sigma(\mathbf{W}_{l-1}f_{l-1}(\mathbf{x})+\mathbf{v}_{l-1}),\quad f_{1}(\mathbf{x})=\sigma(\mathbf{W}_{0}\mathbf{x}+\mathbf{v}_{0}).$ Here, $\mathbf{W}_{l}\in\mathbb{R}^{p_{l+1}\times p_{l}}$ and vectors $\mathbf{v}_{l}\in\mathbb{R}^{p_{l+1}}$ ($l\in[L]$) are the parameters of this DNN, and $\sigma:\mathbb{R}\to\mathbb{R}$ is the activation function, applied element-wise to vectors. A commonly used activation function is the rectified linear unit (ReLU) function (Nair & Hinton, 2010), or $\sigma(x)=\max(x,0).$ We focus on ReLU due to its piecewise linearity and projection properties, with potential extensions to other activation functions as discussed in Section 6.
Overparameterized DNNs can memorize noise and overfit (Cao et al., 2022), so sparsity is used to control complexity. We focus on the following class of $s$-sparse and $F$-bounded networks for our theoretical analysis:
@@ 2.2 Generalized Nonparametric Regression
Consider outcomes from the exponential family, including commonly used distributions such as normal, Bernoulli, Poisson, and binomial. The density (or mass) function for a random variable $y$ in the (single-parameter) exponential family is $p(y|\theta)=h(y)\exp\left\{\theta y-\psi(\theta)\right\},$ where $\theta$ is the canonical parameter, $0<h(y)<\infty$ ensures normalization, and $\psi(\theta)$ is convex and smooth, satisfying $\mathbb{E}(y|\theta)=\psi^{\prime}(\theta)$, where $\psi^{\prime}(\cdot)$ denotes the derivative of $\psi(\cdot)$ and the inverse function of $\psi^{\prime}(\cdot)$ is known as the link function (Brown, 1986; Fei & Li, 2021).
@@
where $f_{0}:\mathbb{R}^{p}\to\mathbb{R}$ is a nonparametric function with $f_{0}(\mathbf{x})$ representing the mean parameter of the response variable $y$ given $\mathbf{x}$. Model (2.3) is referred to as a Generalized Nonparametric Regression Model (GNRM), where the outcomes follow a distribution from the exponential family. With an independent test point $(\mathbf{x}_{{\mathop{\mathrm{new}}}},y_{{\mathop{\mathrm{new}}}})$ and that $y_{{\mathop{\mathrm{new}}}}$, given $\mathbf{x}_{{\mathop{\mathrm{new}}}}$, follows (2.3), we aim to estimate $\mathbb{E}(y_{{\mathop{\mathrm{new}}}}|\mathbf{x}_{{\mathop{\mathrm{new}}}})$ which is $\psi^{\prime}(f_{0}(\mathbf{x}_{{\mathop{\mathrm{new}}}}))$, Consequently, it is of interest to estimate $f_{0}$. By the universal approximation theorem (Hornik et al., 1989), DNNs are one of suitable candidates for estimating $f_{0}$. We consider an optimal DNN estimator, $\widehat{f}_{n}^{{\mathop{\mathrm{opt}}}}$, which minimizes the negative log likelihood function (2.3) over the class of DNNs defined in (2.2), that is,
$\widehat{f}_{n}^{{\mathop{\mathrm{opt}}}}=\operatorname{argmin}_{f\in\mathcal{F}(L,\mathbf{p},s,F)}\frac{1}{n}\sum_{i=1}^{n}\{-y_{i}f(\mathbf{x}_{i})+\psi(f(\mathbf{x}_{i}))\}.$ (2.4)
@@
Nonparametric Gaussian Regression: Given $\mathbf{x}$, we assume that the response variable follows a Gaussian distribution with mean $f_{0}(\mathbf{x})$ and variance $\sigma^{2}$. In this setting, $\sigma^{2}$ does not affect the estimation of the mean function $f_{0}(\mathbf{x})$, which can be seen by examining the likelihood and optimization. The likelihood function for a nonparametric Gaussian regression model is given by $p(y|\mathbf{x})=\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(y-f_{0}(\mathbf{x}))^{2}}{2\sigma^{2}}}$. An application of (2.4) to this setting yields:
$\widehat{f}_{n}^{{\mathop{\mathrm{opt}}}}$ $=$ $\operatorname{argmin}_{f\in\mathcal{F}(L,\mathbf{p},s,F)}\frac{1}{n}\sum_{i=1}^{n}\{-y_{i}f(\mathbf{x}_{i})+\frac{1}{2}f^{2}(\mathbf{x}_{i})\}$
@@
corresponding to mean squared error minimization as in Schmidt-Hieber (2020).
Nonparametric Logistic Regression: To model binary outcomes $y\in\{0,1\}$, we set $h(y)=1$, $\psi(\theta)=\log(1+e^{\theta})$ in (2.3). The estimator (2.4) in this setting can be written as:
$\widehat{f}_{n}^{{\mathop{\mathrm{opt}}}}=\operatorname{argmin}_{f\in\mathcal{F}(L,\mathbf{p},s,F)}\frac{1}{n}\sum_{i=1}^{n}(-y_{i}f(\mathbf{x}_{i})+\log[1+\exp\{f(\mathbf{x}_{i})\}]).$
@@
$\widehat{f}_{n}^{{\mathop{\mathrm{opt}}}}=\operatorname{argmin}_{f\in\mathcal{F}(L,\mathbf{p},s,F)}\frac{1}{n}\sum_{i=1}^{n}[\exp\{f(\mathbf{x}_{i})\}-y_{i}f(\mathbf{x}_{i})].$
@@
Nonparametric Binomial Regression: Given covariates $\mathbf{x}$, the response variable $y$ follows a Binomial distribution with a fixed number of trials $n_{\text{trial}}$ and the success probability $p(\mathbf{x})=\frac{1}{1+e^{-f_{0}(\mathbf{x})}}$. This corresponds to $h(y)=\binom{n_{\text{trial}}}{y}$ and $\psi(\theta)=n_{\text{trial}}\log(1+e^{\theta})$. In this setting, the estimator (2.4) is specified as